# Shahriar Noroozizadeh

### v1

@@ 3.2 Supervisory pressure does not explain geometric memory
Orthogonally, one could attempt to attribute the global geometry to the “global” supervision itself:
(Global supervision may explain global memory) A global geometry arises over local associations since the training involves a (global) path-finding objective.
@@
This hypothesis however is already weakened by design in the path-star task. First, path-finding supervision is not provided for unseen paths, paths which nevertheless exhibit a global geometry. Next, in the hardest-token-only task, path-finding supervision is applied only on the end-points of the path; yet a geometry materializes on intermediate nodes (see Fig.˜16(b)). In fact, we give an even cleaner refutation of this hypothesis by analyzing models trained purely on local supervision (i.e., on edge memorization):
@@
A global geometry emerges even in locally-supervised models as seen in the embeddings of tiny graphs of various architectures in Fig.˜1 and §C.3 and in the heatmaps for the large path-star graph (Figs.˜6(c) and 9(c)). Additionally, such a locally-supervised model can be subsequently finetuned purely on the hardest-token task and achieve high test accuracy on path-finding (§D.3).
To summarize this section, what we demonstrate is a geometric representation that arises in a memorization task. This geometry can be thought of as the crux of broader geometric phenomena in language modeling, here emerging without any of the typical statistical redundancies and learning pressures. This in turn isolates a fundamental aspect of neural geometries that cannot be easily explained. A more nuanced argument is necessary, perhaps by relying on other notions of complexity (e.g., flatness or spectral norms) or by directly analyzing the dynamics.
@@ 4 Geometry arises from naturally-occurring spectral bias, without pressures
Setting aside the competition between the two parametric memories, we can still extract a non-trivial question: how does gradient descent synthesize global information from mere local supervision, without various pressures, and what geometry does it produce? To isolate this, we turn to simpler, 1-layer, 1-hop Node2Vec models. These are equivalent to a Transformer, trained only on edge-memorization, with only an embedding and unembedding layer—thus, associative memory is architecturally prohibited. [^4] Node2Vec cannot implement associative memory at least in the form we care about. A more contrived form is possible; see §E.3.
@@
A precise characterization of what embeddings are learned even in such simple models is an open question, but a rich line of work (albeit with key assumptions about various pressures outlined shortly) points to a spectral bias: the learned embeddings often align with the top (non-degenerate) eigenvectors of the negative graph Laplacian. This indeed holds: the Node2Vec embeddings in Fig.˜1 right column matches the top eigenvectors—called the Fiedler vector(s)—in Fig.˜7 left column. This leads us to a few key insights. First, these eigenvectors happen to be the very source of global geometries. But secondly, these Node2Vec geometries turn out to be more well-organized than what the Transformer exhibited (in Fig.˜1 middle column). Thus, we conjecture a similar but—in hindsight—somewhat “adulterated” spectral geometry in Transformers:
@@
(Spectral bias) A Transformer memorizes with a global geometry due to spectral biases; but there is significant headroom in the quality of geometry, likely because the representation is adulterated with local associative memory.
@@
A natural next question is to wonder where the spectral bias stems from. Going back to Node2Vec theories (see §5.6), the literature suggests that similar models rely on the top eigenvectors due to the explicit pressure of a bottleneck [84, 59, 149, 68] or explicit regularization [75] or an explicitly multi-hop supervision [124]. Our setting defies all these assumptions.
A further discrepancy in existing analyses is that they are in non-cross-entropy-loss settings with simpler losses. In these systems, the dynamics simplify nicely yielding a closed-form solution for the inner products of the embeddings. However, we use the cross-entropy loss (to be faithful to how sequence models are trained), under which an expression for the inner product is evasive. The dynamics here may behave in one of many complex ways: it may simply diverge, or it may converge in direction (like in logistic regression [146]); the converged direction in turn, may be degenerate or not. Our empirical analysis points to a special dynamic: the system does converge to a meaningful zero-gradient solution, working its way towards a neat two-fold property, while exhibiting a spectral bias naturally:
@@
(How spectral bias emerges without typical pressures) In a $1$-layer, 11-hop Node2Vec model with the embedding ${\bf{V}}\in\mathbb{R}^{n\times m}$ of $n$ nodes, with the dynamics denoted as $\dot{{\bf{V}}}(t)=\eta{\bf{C}}(t){\bf{V}}(t)$ (where ${\bf{C}}(t)\in\mathbb{R}^{n\times n}$ is a time-dependent co-efficient matrix), the converged solution for our small graphs is such that (a) the columns of embedding matrix $\bf{V}$} span the graph’s Fiedler-like vectors, and (b) the co-efficient matrix $\bf{C}$} has those same vectors in its null space (Fig.˜7). Crucially, this dynamic does not need a low rank constraint; the embedding size $m$ can be larger than the graph.
@@
Figure 7: Spectral geometry arises in Node2Vec without low-rank pressure (Section˜4) for tiny Path-Star, Grid, Cycle, and Irregular Graphs (top to bottom). (a) The Fiedler-like vectors of the graphs encode global structure; this structure mirrors the Node2Vec embeddings shown in Fig.˜1. (b; left) The evolution of eigenvector projections during training. (middle) The embedding matrix $\bf{V}$} contracts into space spanned by the Fiedler-like eigenvectors, evidenced by the projection norm ${\bf{V}}^{T}{\bf{e}}_{i}||_{2}$ converging to a stable, non-zero value; projections of other eigenvectors diminish towards zero. (b; right) Concurrently, the null space of the co-efficient matrix $\bf{C}$} subsumes the the Fiedler-like eigenvectors, in that the norm ${\bf{C}}{\bf{e}}_{i}||_{2}$ converges to 0. This spectral bias arises without a low dimensional constraint assumed in literature (embedding size $m=100$, much larger than nodes in graph).
@@
We provide in §F an empirically-informed intuition for a “self-stabilizing” dynamic that gradually filters out lower eigenvectors; we leave open a formal analysis. Admittedly, neither is this analysis on a deep sequence model, nor does it divulge anything about the competition between associative and geometric memories. What it does is make progress on an open question for a simpler model. It also gives us an idea of how and what global information can arise naturally out of local supervision, devoid of any supervisory, bottleneck-driven, or regularizing pressure.

### v2

@@ 3.1 Supervisory pressure does not explain geometric memory
Geometric representations arise in various tasks with reasoning objectives, from arithmetic [100, 120, 117, 59] to path-finding [85] to two-hop reasoning [174, 43, 189, 74]. Reasoning supervision has also been shown to act as a regularizer that elicits useful structures in the model’s storage [74]. This trail of findings overwhelmingly directs us to the following guess for why we see a geometry:
(Pressure from global supervision) Perhaps, the model memorizes via a global geometry rather than via local associations due to the supervisory pressure of a multi-hop reasoning objective.
@@
This hypothesis however is weakened by the distinctive design of the path-star task. First, the test paths are untouched by any path-finding supervision, yet these paths exhibit a global geometry. Next, in the hardest-token-only task, path-finding supervision is applied only on the path end-points; yet a geometry materializes on intermediate nodes (see Fig.˜25(b)). As an even cleaner refutation of this hypothesis, we ablate the reasoning supervision, and train purely on local supervision i.e., on edge memorization:
@@
(Geometry arises even from memorizing local, atomic co-occurrences) A global geometry emerges even in locally-supervised models for various architectures, as seen in the heatmaps for the large path-star graph (Figs.˜6(c) and 15(c)). A locally-supervised model can be subsequently finetuned purely on the hardest-token task and achieve high test accuracy on unseen paths (§D.3). We also note that the geometries seen on the small graphs (Fig.˜1; §C.3) are for various architectures trained only on edge-memorization.
From this, we isolate the following insight: geometries can arise instead of an associative lookup, even when memorizing atomic co-occurrences devoid of any pressure to reason or to find high-level, multi-hop structures in data.
3.2 Capacity or optimizer pressures do not explain geometric memory.
We may seek yet other explanations out of canonical intuitions in theories of generalization. In regression or classification, conventional wisdom was that models learn representations rather than a lookup when there is an explicit capacity pressure e.g., the architecture or the regularizers forbid a lookup. In modern theories since the popularization of the generalization puzzle [122, 195], renewed wisdom has been that this pressure is implicit: even if the model can express a lookup, gradient descent implicitly prefers solutions that are easier-to-find or are more succinct, thus leading to useful representations. It may seem that one of these theories, when imported to the world of sequence models, can tell us why a geometric representation arises over a naive lookup of atomic associations.
We scrutinize (straightforward versions of) both these traditional and modern theories as plausible explanations for the geometry, and we refute them one by one. We begin with the traditional notion of explicit capacity pressures:
(Explicit capacity pressures) Perhaps, the model memorizes geometrically because a lookup of local associations is impossible to represent. This restriction is either due to the parameter count or the design of the architecture, especially the embedding bottleneck.
This logic can be dismantled as follows. First, we note that our architectures can express the associative model. Positive theoretical results in Nichani et al. [124] roughly prove that, with $m$ (frozen) embedding dimensions, and at least $m^{2}$ free parameters, an MLP layer can store $m^{2}$ many associations. Our models have $m\approx 400$ which means our large graphs can be stored associatively in our models.
But maybe, there are yet other explicit pressures—such as normalizations and weight decay—that discourage associative storage. As a definitive refutation of such pressures, we make an empirical demonstration: there are experimental settings (with sufficient capacity and no bottleneck pressure) where geometric memory naturally arises; and in those very settings, an associative memory can be learned with the embeddings frozen and all else unchanged. Thus, there is demonstrably no explicit capacity pressure restricting an associative memory—and yet, a geometric one naturally emerges.
(Associative memory can be artificially learned with our architectures) On various sequence models (Transformers, Mamba, neural networks) and various tiny graphs in Fig.˜1 and §C.3, a geometric memory arises naturally even though the same setup fits the data with frozen embeddings. [^3] This fit is purely associative since only one trainable layer exists in these architectures (see §B.2.2).
Thus, explicit capacity pressures are not required to force a geometric memory in place. Maybe then the optimizer has a role to play. For instance, it is known that in classification settings, gradient descent takes longer to discover a lookup of artificially-randomized labels than natural labels [12]; it is also well-known that optimization tends to fit the data in progressively more complex ways [140, 184, 82, 12, 146, 19] and so a lookup would take relatively long to discover than other solutions. A similar effect can be proposed here:
(Optimizational pressure) Perhaps, the model memorizes geometrically because an associative storage is harder (e.g., takes longer) to find by gradient descent.
Refuting this, we argue that in sequence modeling, an associative lookup is significantly easier to find than the geometric one. In fact, during training, models typically first store the data associatively before storing geometrically. First, this is already (indirectly) evidenced at various points in literature (although not interpreted this way). In the graph tasks of [189, 174], the Transformer fits atomic facts well before it “groks” to learn circuits/representations that aid two-hop reasoning; in natural language [96] models memorize n-gram statistics well before representations become more organized, giving way to generalization.
A more precise and vivid claim can be made here: Nichani et al. [124] observe that the associative lookup table $\sum_{(u,v)\in E}{\bf{\Phi}}(u)\otimes{\bf{\Phi}}(v)$ (from Section˜2.3) can be found by one step of gradient descent given a reasonably large embedding size on the correlation loss, $\sum_{(u,v)\in E}f(u)[v]$. [^4] This “one-step-lookup-table” phenomenon is distinctive of sequence modeling but not of a typical classification/regression task: for the lookup to be learned easily, we want the inputs near-orthogonal. This is true by design of the initial token embeddings of any sequence modeling task, but not necessarily true of the input distribution in classification/regression. We empirically verify this for the cross-entropy loss on our tiny graphs: associative memory forms in $2$ steps (Fig.˜7); whereas a geometric memory requires $100$ steps (Fig.˜8).
Our intuition for this “associative first, geometry later” dynamics is that an associative memory merely stores the adjacency matrix $\bf{A}$} that is surficially available in the data; a geometric memory requires unearthing implicit information from the data, which requires more iterations—as we will see in §4, this computation corresponds to low-rank factorizing the adjacency matrix, or equivalently, computing a multi-hop adjacency matrix $\bf{A}$$^{\ell}$.
In summary, an associative memory is easier to find than the geometric one, and yet the model (eventually) organizes its memory geometrically:
(Associative memory is easier to find with gradient descent) Given sufficient embedding dimensionality and capacity, the associative storage can be discovered by full-batch gradient descent in as little as two steps on all our tiny graphs (Fig.˜7) while a geometric memory takes much longer to form (Fig.˜8). Similarly, we interpret the earlier findings of [189, 174, 96] as the data being fit associatively early during training, while a geometry (and therefore, reasoning) arising only much later.
Yet another plausible explanation for the geometry can be derived from the core of modern intuition about generalization, thanks to [122, 195]: gradient descent implicitly finds succinct fits of the data even without being explicitly told to do so. So, even if gradient descent is quick to discover associative memory, over time, perhaps it discovers a geometric organization as it gradually “compresses” the data. Indeed, such compressive or complexity-reducing effects have been brought up in recent analyses of geometry in language [96] and reasoning [174]. We translate this into our terminology:
(Implicit capacity pressure) Perhaps, the model memorizes geometrically (over the course of training) because gradient descent converges to a more succinct fit of the data.
To refute this, we refute the underlying premise itself: counterintuitively, a geometric storage is not necessarily more succinct than an associative storage. To see why, we first clarify that unlike a typical classification task, or even a typical natural language task, our setting must be recognized as a memorization task (in the sense of Feldman [42]): the existence of an edge cannot be statistically surmised from the rest of the training set. Thus, we will show, for certain graphs, straightforward notions of succinctness do not break the tie between the two ways of representing the data.
Concretely, in generalization theories, the complexity of a lookup table is larger than the succinct one by a factor that scales polynomially with the number of datapoints. In contrast, in the path-star or cycle memorization task, this gap (measured through bit or norm complexity) can vanish:
(Geometric and associative memory are roughly equally succinct) For certain graphs, the complexity of the local associative memory both in terms of bits and $\ell_{2}$ norms is either equal to or at most twice those of the global geometric memory. [^5] Equality is achieved if there is no weight-tying or if only forward edges are stored. [^6] Note that a geometry appears even when storing only one direction of the edges in the smaller graphs where both forms of storage must be equally succinct; see §D.1.2.Informal proof and more explanations in §E.2.
Thus, while the familiar geometries of word embedding tasks [40, 130, 112] arise where a lookup is clearly more verbose, similar geometries arise even in tasks where redundancies are absent and a lookup may be just as succinct.
3.3 The memorization puzzle
The above arguments isolate a fundamental aspect of well-known neural geometries that cannot be easily explained by well-known learning pressures. Geometric representations emerge even in an atomic fact-memorization task that lacks statistical redundancies. Attempts to understand this through familiar intuitions from classification/regression tasks—from the lens of ease-of-optimization and succinctness—seem to fail us in the land of sequence models. [^7] Intuitively, this is because the geometric representations are of a fundamentally different nature from the well-studied feature representations in classification/regression tasks (see Remark 3). This gives rise to what we call the memorization puzzle:
Open Question 1: The Memorization Puzzle in Sequence Modeling Why does a gradient-descent-trained sequence model memorize atomic co-occurrences geometrically, rather than associatively, even in cases where the associative storage is expressible, easier-to-find and just as succinct as the geometric storage?
Answering this requires a more thoughtful argument, perhaps by borrowing other notions of complexity (e.g., flatness [84] or spectral norms [123] or something else not considered so far). One may also consider some other well-understood effects of depth and factorization in neural networks [74, 11]; although, we cast doubt even on these effects; see Remark˜1.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 7: Associative memory can be discovered quickly by gradient descent, given a sufficiently wide model, and a sufficiently large learning rate: For the various tiny graphs described in §C.3, and for our TinyNN model (with frozen embedding/unembedding layers and one wide trainable weight matrix to prevent a geometry from taking over; see §B.2.2), we report memorization over timesteps of training with full-batch gradient descent and constant learning rate of $0.1$. For each vertex, we compute what fraction of its $k$ neighbors appear in the top $k$ next token probabilities; this is averaged across vertices (where $k$ varies across vertices). This value, which quantifies associative memorization, reaches its maximum within 22 steps of gradient descent; compare this to the longer time for geometric memory to form in Fig.˜8. This is evidence for our Section˜3.2: ease-of-discovery does not dictate what memory is formed.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 8: Geometric memorization takes much longer for gradient descent to discover: For the same architecture as in Fig.˜7 but with learnable embeddings, we roughly quantify geometric memorization over time via cosine distance across all pairs of vertices, alongside 2D visualizations. Observe that the geometries emerge after about a $200$ steps or more, compared to the 22 steps required for associative memory in Fig.˜7. This is evidence for Section˜3.2. While this is for learning rate $0.001$ (where the geometries are cleanest), the observation holds for larger learning rates too; please see Fig.˜22.
@@ 4 Geometry arises from naturally-occurring spectral bias, without pressures
Setting aside the competition with the associative memory, we can still ask questions about the geometry: what is the geometry, how does it correspond to global information, and how does it arise without explicit pressures? Recall that we find these geometries not only in models with attention layers, but also in Mamba and in classical deep networks; to focus on the core phenomenon, we study the simplest yet non-trivial multi-layered gradient-descent-trained model: a 2-layer linear network consisting only of an embedding and unembedding layer (weight-tied), trained on the same loss (cross entropy) and the same dataset (of local edges). Equivalently, this is a one-layer, 1-hop Node2Vec model. Here, associative memory is architecturally prohibited, [^8] Node2Vec with weight-tying cannot implement associative memory at least in the form we care about. A more contrived form is possible; see §E.3. so we can pursue a clean analysis of the geometry alone.
4.1 Where does the geometry come from?
@@
A precise characterization of what embeddings are learned even in such simple models is an open question. But a rich line of work—with key assumptions about various pressures outlined shortly—points to a low-rank spectral bias: the learned embeddings often align with the top (non-degenerate) eigenvectors of the negative graph Laplacian (roughly given by ${\bf{A}}-{\bf{D}}$, the difference between the adjacency and the diagonal degree matrix). Our figures corroborate this: the Node2Vec embeddings in Fig.˜1 right column match the top eigenvectors—called the Fiedler vector(s)—in Fig.˜9 left column.
This leads us to a few key insights. First, given the similarity in the geometries of Node2Vec and our deep sequence models (Transformer, Mamba and neural networks), such a spectral bias must indeed be at the heart of all such deep models, regardless of architectural quirks like self-attention and residual connections. However, the Node2Vec geometries turn out to be more well-organized than that which the Transformer exhibited (in Fig.˜1 middle column). Thus, we conjecture a similar spectral geometry in Transformers but—in hindsight—mildly adulterated by the competition with local associative memory.
Conjecture 1.
@@
(Adulterated spectral bias) The geometric memory of a Transformer has significant headroom in the quality of its geometry; we conjecture that this is because a gradient-descent-trained Transformer has spectral bias that is adulterated with some level of associative memory.
4.2 Where does the global information come from?
Another key insight is that the top eigenvectors, which are known to encode global structures in the graph, are the very source of global information that aided our reasoning task. Intuitively, the top eigenvectors of the adjacency matrix $\bf{A}$} are also the directions that become dominant in the $\ell$-hop adjacency matrix $\bf{A}$$^{\ell}$, directions that translate to global information. An alternative perspective is to view geometric memory as having factorized the adjacency matrix $\bf{A}$} while (softly) filtering out the lower eigenvectors; this extra computation has produced new information not surficially available in the adjacency data.
In effect, we view Transformers as approximating the global geometry of a Node2Vec. Although only an approximation, the implications are profound: the Transformer comes with the power to reason over the global geometry, an ability lacking in Node2Vec-style models.
4.3 Where does the low-rank spectral bias come from?
@@
While the geometry stems from a low-rank spectral bias, it is not clear where the bias itself comes from. Going back to Node2Vec theories (see §5.5), the literature suggests that Node2Vec-like models rely on the top eigenvectors under the assumptions of an explicit pressure of a bottleneck [93, 66, 165, 76, 152, 153, 14] or explicit regularization [83, 199] or an explicitly multi-hop supervision [137]. Our setting defies all these assumptions.
We hypothesize that these analyses inadvertently introduce explicit pressures in an attempt to simplify the dynamics e.g., nearly all analyses are on simpler-to-analyze losses such as the squared-error or a negative-sampling loss. On the other hand, the (more complex-to-analyze) cross-entropy loss is known to provide implicit pressures in certain other settings, such as in classification where it induces margin-maximization [162]; it is also the loss of choice in modern language models. Thus, we undertake an analysis of this loss, without assuming any explicit pressures. We emphasize that closed-form expressions here have remained an open question.
A priori, one would expect the unregularized cross-entropy-based Node2Vec system to behave in one of few ways: the system may diverge (to drive the probabilities to 0 or $1$), or may converge but only in direction (like in logistic regression [162]); the converged direction in turn, may be degenerate (e.g., node representations collapse to a single value) or not. Through empirical analysis, we suggest that none of these are true. Instead, the system appears to converge to a zero-gradient solution. It does so by elegantly achieving two properties, one of which is a low-rank spectral bias:
@@
(How spectral bias emerges without typical pressures) In a $1$-layer, 11-hop Node2Vec model with the embedding ${\bf{V}}\in\mathbb{R}^{n\times m}$ of $n$ nodes, with the dynamics denoted as $\dot{{\bf{V}}}(t)=\eta{\bf{C}}(t){\bf{V}}(t)$ (where ${\bf{C}}(t)\in\mathbb{R}^{n\times n}$ is a time-dependent co-efficient matrix), the converged solution for our small graphs is such that (a) the columns of embedding matrix $\bf{V}$} span the graph’s Fiedler-like vectors [^9] We use “Fiedler-like” to denote not just the second-top eigenvectors of the negative Laplacian matrix, but also some of the subsequent top eigenvectors., and (b) the co-efficient matrix $\bf{C}$} has those same vectors in its null space (Fig.˜9). Crucially, this dynamic does not need a low rank constraint; the embedding size $m$ can be larger than the graph.
@@
Figure 9: Spectral geometry arises in Node2Vec without low-rank pressure (Section˜4.3) for tiny Path-Star, Grid, Cycle, and Irregular Graphs (top to bottom). (a) The Fiedler-like vectors of the graphs encode global structure; this structure mirrors the Node2Vec embeddings shown in Fig.˜1. (b; left) The evolution of eigenvector projections during training. (middle) The embedding matrix $\bf{V}$} contracts into space spanned by the Fiedler-like eigenvectors, evidenced by the projection norm ${\bf{V}}^{T}{\bf{e}}_{i}||_{2}$ converging to a stable, non-zero value; projections of other eigenvectors diminish towards zero. (b; right) Concurrently, the null space of the co-efficient matrix $\bf{C}$} subsumes the Fiedler-like eigenvectors, in that the norm ${\bf{C}}{\bf{e}}_{i}||_{2}$ converges to 0. This spectral bias arises without a low dimensional constraint assumed in literature (embedding size $m=100$, much larger than nodes in graph).
Why do the dynamics play out this way? We provide in §F an empirically-informed intuition (and a partial proof) for a “self-stabilizing” dynamic. The system gradually filters out lower eigenvectors from the embedding matrix $\bf{V}$}, while increasing their presence in the co-efficient matrix $\bf{C}$}, zeroing out the gradients ${\bf{C}}{\bf{V}}$ over time. We leave open a formal analysis of convergence, its rate, and whether this can be characterized into a form of max-margin-style bias.
@@
We emphasize that this two-layer analysis does not divulge anything about the competition between associative and geometric memories. That requires a study of deeper models, where the dynamics of different spectral components begin cross-interfering. What this two-layer analysis does is make progress on an open question for the simplest model possible. It also gives us an idea of how and what global information can arise naturally out of local supervision, devoid of any supervisory, bottleneck-driven, or regularizing pressure. These spectral dynamics must be central to any gradient-descent-trained deep sequence model.
With the above insights from Node2Vec, we revisit our original deep sequence models where the geometric memory is in competition with, and adulterated by, associative memory (as noted in §4.1); here, we pose the following concrete question(s) in quantifying this competition:
Open Question 2: Characterizing and Controlling the Geometric Bias (i) To what precise extent does a deep sequence model memorize different parts of a graph dataset geometrically vs. associatively? (ii) What local or global notions of graph complexity—such as the graph spectrum, connectivity, node & vertex count and degree—and what aspects of optimization—such as the architecture, learning rate, training time —control this bias? (iii) What interventions in the data, architecture or optimization can be made to make memorization more geometric or more associative as desired in practice?
4.4 The importance of weight-tying
Our Node2Vec model shares weights between the two layers, reflecting the de facto practice of tying unembedding and embedding matrices in language models. This convention coincidentally plays a unique role in memory: without it, the geometries do not seem visible, at least on the face of it (see Fig.˜10 for Node2Vec and Fig.˜11 for Transformer, and the corresponding multi-hop cosine-similarities in Fig.˜33 and Fig.˜34).
We give an explanation for this. A weight-untied two-layer model must factorize the adjacency matrix not as a eigenvalue decomposition (EVD) but as a singular value decomposition (SVD)—this intuitively follows from analyses of weight-untied Word2Vec-style models [199, 83, 152]. The eigenvalues of the adjacency matrix range both positive and negative; the SVD “folds up” the negative eigendirections—all strong eigendirections, both the most positive and most negative, become top singular directions (see [197] for relevant facts on spectral graph theory). The most negative eigendirections however assign zig-zagging values to adjacent nodes (e.g., +1+1 and 1-1 as we walk along a path). This manifests as the unruly figures in Figs.˜10 and 11.
This raises various questions. First, although visually unruly, these zig-zagging eigendirections are highly-structured: they approximate a bipartite cut of the graph. Thus, they correspond to a form of geometric memory distinct from the “global” one, best termed as a bipartite geometric memory. Next, while the “global” geometry endows the model with navigational powers (as witnessed in §2), what benefits does the bipartite memory bring to a model’s reasoning abilities?
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 10: Embeddings of weight-untied Node2Vec do not show a clean geometry in the top directions. Corresponding multi-hop cosine similarities are in Fig.˜33.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 11: Embeddings of a weight-untied Transformer do not show a clean geometry in the top directions. Corresponding multi-hop cosine similarities are in Fig.˜34.